# A review of Bayesian sensor-based estimation and uncertainty quantification of aerodynamic flows

### v2

@@ Abstract
Many applications in aerodynamics depend on the use of sensors to estimate the evolving state of the flow. In particular, a wide variety of traditional and learning-based strategies for closed-loop control rely on some knowledge of the aerodynamic state in order to decide on actions. This estimation task is inherently accompanied by uncertainty due to the noisy measurements of sensors or the non-uniqueness of the underlying mapping, and knowledge of this uncertainty can be as important for decision-making as that of the state itself. The tracking of uncertainty is challenged by the often-nonlinear relationship between the sensor measurements and the flow state. For example, a collection of passing vortices leaves a footprint in wall pressure sensors that depends nonlinearly on the strengths and positions of the vortices. In this paper, we will review the recent body of work on flow estimation. We will discuss the basic tools of probability, including sampling and estimation, in the powerful setting of Bayesian inference and demonstrate these tools in static flow estimation examples. We will then proceed to unsteady examples and illustrate the application of sequential estimation, and particularly, the ensemble Kalman filter. Finally, we will discuss uncertainty quantification in neural network approximations of the mappings between sensor measurements and flow states. Recent aerodynamic applications of neural networks have shown that the flow state can be encoded into a very low-dimensional latent space, and we will discuss the implications of this encoding on uncertainty.
keywords:
Flow estimation , Bayesian inversion , data assimilation , uncertainty quantification , unsteady aerodynamics
journal: Progress in Aerospace Sciences
[…]
organization=University of California, Los Angeles,addressline=420 Westwood Plaza, Box 951597, city=Los Angeles, postcode=90095, state=CA, country=USA
@@ 1 Introduction
The focus of this paper is on addressing problems in which we seek to estimate the state of a fluid flow from a set of sensor measurements. Though we will not address the purpose of this flow estimation, it ostensibly fits in with some larger task of flow or flight control: We would like to act to achieve some objective (maintain steady forces, perform agile maneuvers) in the presence of atmospheric disturbances (i.e., gusts) [1]. The signature of these disturbances can be found in the sensor data, and we would like to use these data to form a more complete picture of the flow.
@@ 2 Static inference of flows
In this section we will address the problem of inverting (1) in a probabilistic sense. That is, we will think of the state $\mathsf{x}$ and observations $\mathsf{y}$ as continuous random variables, and ask the question: given $\mathsf{y}$, what is the most likely $\mathsf{x}$? To address this question, we will review several concepts and tools from probability theory. This review is meant to be cursory and illustrative, and assumes some basic knowledge of probability. For more detailed coverage, the reader is referred to any of several books on the subject. We particularly recommend the books by Bishop [2] and Asch et al. [3], both of which have one or more chapters devoted to basics of Bayesian probability. We also highly recommend the review by Stuart [4].
@@
which sensibly puts more weight on more likely states. Note that the function $F(\mathsf{x})$ could be just $\mathsf{x}$ itself, in which case (4) returns the expected state (the conditional mean). One could also, for example, compute the expected state covariance to assess our uncertainty of the expected state. Equations (43)–(47) in A summarize the most important quantities.
@@ 2.1 The Bayesian view of inversion
This theorem is simple but provides a very powerful way of framing the inversion problem in a probabilistic manner [4]: We start with some basic knowledge of the distribution of possible states $\pi_{0}(\mathsf{x})$, called the prior distribution. Our objective is to tighten our knowledge of $\mathsf{x}$ by obtaining the posterior distribution, $\pi(\mathsf{x}|\mathsf{y})$, i.e., the distribution of states conditioned on a measurement. The connection between the posterior and prior is through the likelihood. The denominator, $\pi(\mathsf{y})$, is a marginal distribution of the possible measurement values integrated across all possible states. But Bayes’ theorem is evaluated at a particular value of the observation vector—the true measurement vector, $\mathsf{y}^{\ast}$—and remains only a function of $\mathsf{x}$, so $\pi(\mathsf{y}^{\ast})$ is simply a uniform scaling factor. Often its value is not important, since we only care about comparing the probabilities of two states, that is,
@@
Often it is safe to assume that this noise is Gaussian with zero mean: that is, $\varepsilon\sim\mathcal{N}(0,\Sigma_{\mathcal{E}})$, where the covariance $\Sigma_{\mathcal{E}}\in\mathbb{R}^{d\times d}$ is diagonal (i.e., the sensors’ noise is independent of one another). Equation (8) then implies that the measurements are also Gaussian distributed, about $h(\mathsf{x})$, i.e., $L(\mathsf{y}|\mathsf{x})=\mathcal{N}(\mathsf{y}|h(\mathsf{x}),\Sigma_{\mathcal{E}})$, indicating that we trust our predicted observation only to within the measurement noise. Multiplying this likelihood with the prior gives us access to the posterior distribution.
@@
Figure 1: Examples of the construction of the posterior covariance for observation model $h(\mathsf{x})=2\mathsf{x}$ and a Gaussian prior with varying degrees of noise and a true measurement $\mathsf{y}^{\ast}=1.4$. The gray region depicts the covariance ellipse of the joint distribution. (a) $\sigma_{\mathcal{E}}=0.1$, (b) $\sigma_{\mathcal{E}}=0.5$
Bayes theorem has been used by numerous investigators to frame ill-posed inverse problems in a probabilistic setting. As an example in the context of fluid dynamics, Kontogiannis et al. [5] have recently used a Bayesian framework to assimilate noisy images of a flow field (from magnetic resonance velocimetry) to reconstruct the flow through a channel and infer the channel’s geometry. The prior enables them to restrict the estimated flow to solutions of the Navier–Stokes equations. The likelihood expresses the error between the noisy velocity measurements and the velocity computed on a mesh and interpolated to the same locations as the measurements.
We can illustrate the idea of the Bayesian approach to inversion with a simple example. Suppose that $\mathsf{x}$ and $\mathsf{y}$ are both scalar-valued, and the observation operator is linear, $h(\mathsf{x})=2\mathsf{x}$. We will assume the prior is Gaussian with mean $\mu_{\mathsf{x}}=0.5$ and standard deviation $\sigma_{\mathsf{x}}=0.3$. Figure 1 depicts the resulting posterior distribution for two different levels of measurement noise when the true measurement is $\mathsf{y}^{\ast}=1.4$. The joint distribution, shown as a gray covariance ellipse, is assembled from the product of the prior (which sets the ellipse’s horizontal extent) and the likelihood (the vertical width, determined by the noise distribution, shown as a green arrow). The posterior mean then lies at the center of the intersection between the true measurement (a red horizontal line) and the ellipse, and the posterior covariance—the uncertainty of our new estimate—is set by the maximum horizontal width of the covariance ellipse (the red arrow). In one case, the measurement noise is low ($\sigma_{\mathcal{E}}=0.1$), and the resulting posterior is much narrower than the prior and centered at a mean of nearly $0.7$, the value we would get if we inverted without noise. But with larger measurement noise ($\sigma_{\mathcal{E}}=0.4$), the true measurement is trusted less, so the posterior relies more on the prior, lying just to the right of it and only slightly narrower.
@@ 2.2 Linear observation operators and Gaussian distributions
This simple scalar example with a linear observation operator and Gaussian noise and prior distributions extends nicely to multidimensional problems, in which the observation operator is a matrix, $h(\mathsf{x})=H\mathsf{x}$. This extension follows because products of multivariate Gaussians are also Gaussian [2], so, if the prior is $\pi_{0}(\mathsf{x})=\mathcal{N}(\mathsf{x}|\mu_{\mathsf{x}},\Sigma_{\mathsf{x}})$, then the posterior is also Gaussian, $\pi(\mathsf{x}|\mathsf{y}^{\ast})=\mathcal{N}(\mathsf{x}|\mu_{\mathsf{x}|\mathsf{y}^{\ast}},\Sigma_{\mathsf{x}|\mathsf{y}^{\ast}})$, where the posterior mean and covariance are (see A)
@@
When the measurement noise goes to zero uniformly (or the prior uncertainty is large), the Kalman gain in (12) reduces to the pseudo-inverse, $(H^{T}H)^{-1}H^{T}$, and the posterior covariance in (10) to $(H^{T}\Sigma_{\mathcal{E}}^{-1}H)^{-1}$, indicating full trust in the measurements over the prior. In contrast, when the measurements are very noisy (and untrustworthy), then the Kalman gain itself shrinks to zero, keeping the posterior near the prior.
A simple fluid dynamics example
In multi-dimensional problems, the Kalman gain can be unpacked to reveal a lot of interesting detail about how the measurements inform the inference of the state. We will use a very simple fluid dynamics task to demonstrate this. Suppose we have a U-tube manometer for measuring pressure in a test section relative to the ambient, $\Delta p=p-p_{\infty}$. However, we wish to use it to infer the absolute pressures, $p$ and $p_{\infty}$. Clearly this is impossible, but it is illustrative to see how this impossible task fails. Our measurement vector is one dimensional, $\mathsf{y}=\Delta p$, our state vector is two dimensional, $\mathsf{x}=[p\,\,\,p_{\infty}]^{T}$, and the observation operator is the $1\times 2$ matrix $H=[1\,\,\,-1]$. We will assume that the ambient pressure has historically been within 100 Pa of 1 atm ($101325$ Pa), and we do not have any further information about the pressure in the test section, so we will set the prior to be Gaussian, with mean and covariance
@@
[…]
@@
The first column of $V$ represents differences in the entries in $\mathsf{x}$, reflecting the manometer’s function. In (16), the corresponding component of $\breve{\mathsf{x}}$ is multiplied by the only non-zero singular value to produce the sole component of $\breve{\mathsf{y}}$, and from that, the predicted measurement $\mathsf{y}$. This first column of $V$ thus spans the informative subspace of $H$. In contrast, the second column of $V$ represents the sum of the entries in $\mathsf{x}$ (i.e., the average of the pressures), and this component is multiplied by zero. In other words, the average of the pressures lies in the null space of $H$, since the manometer has no ability to sense this average; this second column spans the noninformative, or non-observable, subspace of the state space. In general, when the number of sensors is less than the number of states, the observation matrix is rank deficient: $S$ amplifies some of the state components while diminishing or completely removing others. Those that are removed are not observable by the sensors; the remaining ones are observable to varying degrees.
@@
The diagonal $n\times n$ posterior covariance matrix $\Sigma_{\breve{\mathsf{x}}|\mathsf{y}^{\ast}}=\left(I_{n}+S^{T}S\right)^{-1}$ has entries that are less than 1 where the singular values in $S$ are nonzero, and equal to 1 where the singular values vanish. This has the effect of reducing the covariance in (18) only in the observable directions of $V$, and leaving the non-observable directions the same as in the prior. This is easily seen in Figure 2, in which the uncertainty in the observable direction $(1/\sqrt{2},-1/\sqrt{2})$ (associated with pressure differences) is shrunk, but the non-observable direction $(1/\sqrt{2},1/\sqrt{2})$ (associated with average pressures) is untouched.
The decomposed form of the Kalman gain (20) reveals that the first operation that $K$ applies to the measurement innovation in (9) is to scale it by $\Sigma_{\mathcal{E}}^{1/2}$: sensors with larger noise are thereby assigned lower values in this whitened form. The scaled innovation is then rotated into a new coordinate system by $U^{T}$, in which the modal Kalman gain $\breve{K}$ can act upon each component individually. In equation (21), the zero entries in the matrix $S^{T}$ ensure that the non-observable components of the state are not updated by the sensors. In the manometer example, the mean is shifted only in the informative direction $(1/\sqrt{2},-1/\sqrt{2})$.
@@ 2.3 Nonlinear observation operators
which cannot generally be reduced to a quadratic form for $\mathsf{x}$, as we would need for a Gaussian posterior. Thus, we cannot readily apply the tools of the previous section to find the posterior distribution analytically (though, as we will show, these tools remain valuable). We also cannot analytically compute the marginal distribution in the denominator of (6).

### v3

# A practical guide to estimation and uncertainty quantification of aerodynamic flows

@@ Abstract
Many applications in aerodynamics, particularly in closed-loop control, depend on sensors to estimate the evolving state of the flow. This estimation task is inherently accompanied by uncertainty due to the noisy measurements of sensors or the non-uniqueness of the underlying mapping. Knowledge of this uncertainty can be as important for decision-making as that of the state itself. Uncertainty tracking is challenged by the often-nonlinear relationship between the measurements and the flow state. For example, a collection of passing vortices leaves a footprint in wall pressure that depends nonlinearly on the vortices’ strengths and positions. In this paper, we outline recent approaches to flow estimation and illuminate them with worked examples and selected case studies. We review relevant probability tools, including sampling and estimation, in the powerful setting of Bayesian inference and demonstrate these in static flow estimation examples. We then review unsteady examples and illustrate the application of sequential estimation, and particularly, the ensemble Kalman filter. Finally, we discuss uncertainty quantification in neural network approximations of the mappings between sensor measurements and flow states. Recent aerodynamic applications have shown that the flow state can be encoded into a very low-dimensional latent space. We discuss the uncertainty implications of this encoding.
@@ 1 Introduction
The focus of this paper is on addressing problems in which we seek to estimate the state of a fluid flow from a set of sensor measurements. Though we will not address the purpose of this flow estimation, it ostensibly fits in with some larger task of flow or flight control: We would like to act to achieve some objective (maintain steady forces, perform agile maneuvers) in the presence of atmospheric disturbances (i.e., gusts) [jones2022physics]. The signature of these disturbances can be found in the sensor data, and we would like to use these data to form a more complete picture of the flow.
@@
We apologize preemptively that this paper will fall well short of a comprehensive review of the recent work on the vibrant field of flow estimation. Our objective here is to provide an overview and brief tutorial on state-of-the-art tools and to provide selected examples from the literature that illustrate these tools.
@@ 2 Static inference of flows
In this section we will address the problem of inverting (1) in a probabilistic sense. That is, we will think of the state $\mathsf{x}$ and observations $\mathsf{y}$ as continuous random variables, and ask the question: given $\mathsf{y}$, what is the most likely $\mathsf{x}$? To address this question, we will review several concepts and tools from probability theory. This review is meant to be cursory and illustrative, and assumes some basic knowledge of probability. For more detailed coverage, the reader is referred to any of several books on the subject. We particularly recommend the books by Bishop [bishop2006pattern] and Asch et al. [asch2016data], both of which have one or more chapters devoted to basics of Bayesian probability. We also highly recommend the reviews by Stuart [stuart2010inverse] and by Ghattas and Willcox [ghattas2021learning], the latter of which is devoted to addressing the dovetailing tasks of inverse problems and model reduction. We also direct the reader’s attention to the excellent book by Tarantola [tarantola2005inverse], which provides more depth to the probabilistic setting of the inverse problem.
@@
which sensibly puts more weight on more likely states. Note that the function $F(\mathsf{x})$ could be just $\mathsf{x}$ itself, in which case (4) returns the expected state (the conditional mean). One could also, for example, compute the expected state covariance to assess our uncertainty of the expected state. Equations (43)–(47) in Appendix A summarize the most important quantities.
@@ 2.1 The Bayesian view of inversion
This theorem is simple but provides a very powerful way of framing the inversion problem in a probabilistic manner [stuart2010inverse]: We start with some basic knowledge of the distribution of possible states $\pi_{0}(\mathsf{x})$, called the prior distribution. Our objective is to tighten our knowledge of $\mathsf{x}$ by obtaining the posterior distribution, $\pi(\mathsf{x}|\mathsf{y})$, i.e., the distribution of states conditioned on a measurement. The connection between the posterior and prior is through the likelihood. The denominator, $\pi(\mathsf{y})$, is a marginal distribution of the possible measurement values integrated across all possible states. But in using Bayes’ theorem to obtain the posterior distribution, it is evaluated at a particular value of the observation vector—the true measurement vector, $\mathsf{y}^{\ast}$—and remains only a function of $\mathsf{x}$, so $\pi(\mathsf{y}^{\ast})$ is simply a uniform scaling factor. Often its value is not important, since we only care about comparing the probabilities of two states, and we are content to know the following:
@@
Often it is safe to assume that this noise is Gaussian with zero mean: that is, $\varepsilon\sim\mathcal{N}(0,\Sigma_{\mathcal{E}})$, where the covariance $\Sigma_{\mathcal{E}}\in\mathbb{R}^{d\times d}$ is diagonal (i.e., the sensors’ noise is independent of one another). Equation (8) then implies that the measurements are also Gaussian distributed, about $h(\mathsf{x})$, i.e., $L(\mathsf{y}|\mathsf{x})=\mathcal{N}(\mathsf{y}|h(\mathsf{x}),\Sigma_{\mathcal{E}})$, indicating that we trust our predicted observation only to within the measurement noise. Multiplying this likelihood with the prior gives us access to the posterior distribution. [^3] The likelihood’s evaluation at the true measurement could be thought of as a product of two distributions: an exact distribution representing the (perfect) observation model, and a distribution about the true measurement representing sensor noise. This is the view described by Tarantola [tarantola2005inverse].
@@
Figure 1: Construction of the posterior distribution (red) for observation model $h(\mathsf{x})=2\mathsf{x}$ and Gaussian prior (blue) with varying degrees of measurement noise (green) and a true measurement $\mathsf{y}^{\ast}=1.4$. Covariance ellipse of the joint distribution shown in gray. (a) $\sigma_{\mathcal{E}}=0.1$, (b) $\sigma_{\mathcal{E}}=0.5$
Bayes’ theorem has been used by numerous investigators to frame ill-posed inverse problems in a probabilistic setting. As an example in the context of fluid dynamics, Kontogiannis et al. [kontogiannis2022joint] have recently used a Bayesian framework to assimilate noisy images of a flow field (from magnetic resonance velocimetry) to reconstruct the flow through a channel and infer the channel’s geometry. The prior enables them to restrict the estimated flow to solutions of the Navier–Stokes equations. The likelihood expresses the error between the noisy velocity measurements and the velocity computed on a mesh and interpolated to the same locations as the measurements.
We can illustrate the idea of the Bayesian approach to inversion with a simple example. Suppose that $\mathsf{x}$ and $\mathsf{y}$ are both scalar-valued, and the observation operator is linear, $h(\mathsf{x})=2\mathsf{x}$. We will assume the prior is Gaussian with mean $\mu_{\mathsf{x}}=0.5$ and standard deviation $\sigma_{\mathsf{x}}=0.3$. Figure 1 depicts the resulting posterior distribution for two different levels of measurement noise when the true measurement is $\mathsf{y}^{\ast}=1.4$. The joint distribution, shown as a gray covariance ellipse, is assembled from the product of the prior (which sets the ellipse’s horizontal extent) and the likelihood (the vertical width, determined by the noise distribution, shown in green). The posterior mean then lies at the center of the intersection between the true measurement (a red horizontal line) and the ellipse, and the posterior covariance—the uncertainty of our new estimate—is set by the maximum horizontal width of the covariance ellipse (the red arrow). In one case, the measurement noise is low ($\sigma_{\mathcal{E}}=0.1$), and the resulting posterior is much narrower than the prior and centered at a mean of nearly $0.7$, the value we would get if we inverted without noise. But with larger measurement noise ($\sigma_{\mathcal{E}}=0.4$), the true measurement is trusted less, so the posterior relies more on the prior, lying just to the right of it and only slightly narrower.
@@ 2.2 Linear observation operators and Gaussian distributions
This simple scalar example with a linear observation operator and Gaussian noise and prior distributions extends nicely to multidimensional problems, in which the observation operator is a matrix, $h(\mathsf{x})=H\mathsf{x}$. This extension follows because products of multivariate Gaussians are also Gaussian [bishop2006pattern], so, if the prior is $\pi_{0}(\mathsf{x})=\mathcal{N}(\mathsf{x}|\mu_{\mathsf{x}},\Sigma_{\mathsf{x}})$, then the posterior is also Gaussian, $\pi(\mathsf{x}|\mathsf{y}^{\ast})=\mathcal{N}(\mathsf{x}|\mu_{\mathsf{x}|\mathsf{y}^{\ast}},\Sigma_{\mathsf{x}|\mathsf{y}^{\ast}})$, where the posterior mean and covariance are (see A)
@@
When the measurement noise is uniform and goes to zero (or the prior uncertainty is large), the Kalman gain in (12) reduces to the pseudo-inverse, $(H^{T}H)^{-1}H^{T}$, and the posterior covariance in (10) to $(H^{T}\Sigma_{\mathcal{E}}^{-1}H)^{-1}$, indicating full trust in the measurements over the prior. In contrast, when the measurements are very noisy (and untrustworthy), then the Kalman gain itself shrinks to zero, keeping the posterior near the prior.
A simple fluid dynamics example.
In multi-dimensional problems, the Kalman gain can be unpacked to reveal a lot of interesting detail about how the measurements inform the inference of the state. We will use a very simple fluid dynamics task to demonstrate this. Suppose we have a U-tube manometer for measuring pressure in a test section relative to the ambient, $\Delta p=p-p_{\infty}$. However, we wish to use it to infer the absolute pressures, $p$ and $p_{\infty}$. Clearly this is impossible, but it is illustrative to see how this impossible task fails. Our measurement vector is one dimensional, $\mathsf{y}=\Delta p$, our state vector is two dimensional, $\mathsf{x}=[p\,\,\,p_{\infty}]^{T}$, and the observation operator is the $1\times 2$ matrix H=[11]H=[1\,\,\,-1]. We will assume that the ambient pressure has historically been within 100 Pa of 1 atm ($101325$ Pa), and we do not have any further information about the pressure in the test section, so we will set the prior to be Gaussian, with mean and covariance
@@
Figure 2 depicts the posterior mean and covariance that results when we take a manometer measurement of $\mathsf{y}^{\ast}=-30$ Pa. Whereas the prior distribution had a mean of zero and circular covariance because of our identical uncertainty about both $p$ and $p_{\infty}$, the new mean is at $\mu_{\mathsf{x}|\mathsf{y}^{\ast}}=[101310\,\,\,101340]^{T}$, whose components differ by 30-30, and the new covariance ellipse is still as broad as the prior circle in one direction but much narrower in the perpendicular direction. How can we understand this result? We will address this question by decomposing the observation matrix $H$ with a singular value decomposition (SVD); to normalize the state and observation vectors, we will also re-scale the matrix with the noise and prior covariances,
@@
The first column of $V$, the vector $v_{1}$, represents differences in the entries in $\mathsf{x}$, reflecting the manometer’s function. In (16), the corresponding component of $\breve{\mathsf{x}}$ is multiplied by the only non-zero singular value to produce the sole component of $\breve{\mathsf{y}}$, and from that, the predicted measurement $\mathsf{y}$. This first column of $V$ thus spans the informative subspace of $H$. In contrast, the second column of $V$ (vector $v_{2}$) represents the sum of the entries in $\mathsf{x}$ (i.e., the average of the pressures), and this component is multiplied by zero. In other words, the average of the pressures lies in the null space of $H$, since the manometer has no ability to sense this average; this second column spans the noninformative, or non-observable, subspace of the state space. In general, when the number of sensors is less than the number of states, the observation matrix is rank deficient: $S$ amplifies some of the state components while diminishing or completely removing others. Those that are removed are not observable by the sensors; the remaining ones are observable to varying degrees. Examples 2.1 and 2.2 by Stuart [stuart2010inverse] provide nice illustrations of under- and overdetermined systems.
@@
The diagonal $n\times n$ posterior covariance matrix $\Sigma_{\breve{\mathsf{x}}|\mathsf{y}^{\ast}}=\left(I_{n}+S^{T}S\right)^{-1}$ has entries that are less than 1 where the singular values in $S$ are nonzero, and equal to 1 where the singular values vanish. This has the effect of reducing the covariance in (18) only in the observable directions of $V$, and leaving the non-observable directions the same as in the prior. This is easily seen in Figure 2, in which the uncertainty in the observable direction $v_{1}=(-1/\sqrt{2},1/\sqrt{2})$ (associated with pressure differences) is shrunk, but the non-observable direction $v_{2}=(1/\sqrt{2},1/\sqrt{2})$ (associated with average pressures) is untouched.
The decomposed form of the Kalman gain (20) reveals that the first operation that $K$ applies to the measurement innovation in (9) is to scale it by $\Sigma_{\mathcal{E}}^{1/2}$: sensors with larger noise are thereby assigned lower values in this whitened form. The scaled innovation is then rotated into a new coordinate system by $U^{T}$, in which the modal Kalman gain $\breve{K}$ can act upon each component individually. In equation (21), the zero entries in the matrix $S^{T}$ ensure that the non-observable components of the state are not updated by the sensors. In the manometer example, the mean is shifted only in the informative direction $(-1/\sqrt{2},1/\sqrt{2})$.
@@ 2.3 Nonlinear observation operators
which cannot generally be reduced to a quadratic form for $\mathsf{x}$, as we would need for a Gaussian posterior. Thus, we cannot readily apply the tools of the previous section to find the posterior distribution analytically (though, as we will show, these tools remain valuable). We also cannot analytically compute the marginal distribution in the denominator of (6), if we need it.